silvernight269:

If p is 40% of q and q is 120% or r, then what percent of r is p?

silvernight269:

@Hero

TheSmartOne:

Do you know how to translate that into an equation? Hint: is can be replaced with = of can be replaced with * So for example 50 is what percent of 100 Can be translated to 50 = x% * 100 or 50 = x/100 * 100 Since % means over 100

silvernight269:

That doesnt rly help. i got (A+(A x 20%) - (B x 10%) = 432 (A+(A x 20%) is B

silvernight269:

@TheSmartOne

TheSmartOne:

wat.

silvernight269:

Thats the equation I got. Idk if its right or if it wrong.

TheSmartOne:

I'm struggling to understand how you came up with that

silvernight269:

Me too.

silvernight269:

wait nvm was thinking of another question

silvernight269:

My fualt

TheSmartOne:

So let's go step by step If p is 40% of q Can you write that as an equation based on what I said in my first reply in this thread

silvernight269:

p=40% x q

TheSmartOne:

Perfect Now similarly q is 120% of r

silvernight269:

q=120% x r

TheSmartOne:

Good job. Now we have p= 40% * q q= 120% * r And we want to find what percent of r is p So we're trying to find p = x% * r And we're trying to find this x value Now how do we go about this. Do you have any idea? Like maybe we can substitute one equation into the other

silvernight269:

Plug in the 40% x q into p?

TheSmartOne:

So let's look at it a different way You know that if a= b and b = c Then a = c So here we have p = 40% q q = 120% r so we can substitute one of them into the other and get rid of one of the variables. Do you see it?

silvernight269:

40% * 120% * r

TheSmartOne:

Exactly So p = 40% * 120% * r

TheSmartOne:

So let's figure out what 40% * 120% is Percent (%) means over 100 We can just easily convert the percentages into a decimal and then multiply them. And then go back to a percentage So what is 0.4 * 1.20

silvernight269:

0.48

TheSmartOne:

And now what would that be as a percentage

silvernight269:

48%

TheSmartOne:

Bingo

TheSmartOne:

So what percent of r is p?

silvernight269:

great

silvernight269:

48

TheSmartOne:

Perfecto. You got it. Good job :)